If a basement sump pump can empty a 5.0 gallon basin (approximately 19 kg) of water in 4.5 seconds and send the water out of a house through an outlet pipe 2.5 meter above the basin, how many watts must the pump motor minimally be rated at? (Assume frictionless pipes!) How many horse power is that?


1
Expert's answer
2021-03-23T11:10:34-0400

Find the potential energy the water acquires:


"E=mgh."

The work required to lift this mass of water:


"W=\Delta E=E_f-E_i=E-0=E=mgh."

The power is how quickly the work can be done:


"P=\frac Wt=\frac Et=\frac{mgh}{t},\\\space\\
P=\frac{19·9.8·2.5}{4.5}=103\text{ W}."

This is equivalent to 0.138 hp because 1 watt is 0.00134 horse power.